Question 13 Not yet answered Marked out of 10.00 Solve the following initial value problem. dy/dx + 3xy - 2x e^(-x) = 0 (firsPlease circle your final answer!
Answer #1

equa tion is (13) Given n differential dy + 3x²y - zaé -x] die +3ey xe x² dx = 2xe ولے x² T + 3x², y=zxé is in the The equat3 (%) T-FE R) =e 3 ) .. 3 I. ت نع e Then the general soluation is y.(IF) - 5 am). (T-F) dete (6x3, dx +C دیدم saxet علا 2 3د yet = xt tc ای - 3 الحاد) ,ا کرم رہ کرے یا (2) To finds Given initial condition ل = (م) و ههه Now equation o becomes -3 (eº=1 ri al 1000 = I (O+C) looo ICC) (200 C .. c=looo substitute Cralue in equation equation ③ becomes ya Now ex (² +1000)=(x²+1000) et Hence 3 y =(2²+1000) k which is required Soluation Note The option å is Correct answer